A charge of 7.15 mC is placed at each corner of a square 0.450 m on a side.

Determine the magnitude of the force on each charge.

Determine the direction of the force on a charge.

Determine the direction of the force on a charge.

along the line between the charge and the center of the square outward of the center
along the line between the charge and the center of the square toward the center
along the side of the square outward of the other charge that lies on the side
along the side of the square toward the other charge that lies on the side
Answer #1

q2 a &q=%-15 me RF 23 That I qA a qu as shown in the figure, het force Fhet = F2 + √2 F1 = 4 (202) + unto Q² at .92 +5. I. 82